# Diffusion Reward: Learning Rewards via Conditional Video Diffusion

### 21 Dec 2023 v1

@@ 1 Introduction
Figure 1: Overview. (top) We present a reward learning framework in RL using video diffusion models. We perform diffusion processes conditioned on historical frames to estimate conditional entropy as rewards to encourage RL exploration of expert-like behaviors. (bottom) The mean success rate of 10 visual robotic manipulation tasks demonstrates the effectiveness of our Diffusion Reward over 3 runs. Shaded areas are 95% confidence intervals.
@@ 3 Preliminaries
Real VideoGPT 0.526 (±$\pm 0.150$) 18.61 (±$\pm 5.29$) 0.1184 (±$\pm 0.0498$)
VQ-Diffusion 0.955 (±$\pm 0.009$) 33.19 (±$\pm 0.88$) 0.0151 (±$\pm 0.0052$)
Sim. VideoGPT 0.723 (±$\pm 0.086$) 18.70 (±$\pm 2.23$) 0.1175 (±$\pm 0.0513$)
VQ-Diffusion 0.732 (±$\pm 0.085$) 19.69 (±$\pm 2.82$) 0.1078 (±$\pm 0.0615$)
@@ 4.3 Training Details
Figure 6: Main results. Success rate for our method and baselines on 7 gripper manipulation tasks from MetaWorld and 3 dexterous manipulation tasks from Adroit with image observations. Our method achieves prominent performance on all tasks, and significantly outperforms baselines on complex door and hammer tasks. Results are means of 3 runs with 95% confidence intervals (shaded area).
@@ 5.1 Experimental Setup
Figure 7: Success rate curves on 5 unseen MetaWorld tasks. Diffusion Reward could generalize to unseen tasks directly and produce reasonable rewards, largely surpassing other baselines. Results are means of 4 runs with 95% confidence intervals (shaded area).
@@ 5.2 Main Results
We present the learning curves of success rates for each method over two simulation domains in Figure 1 and 6. The results show that merely using sparse environmental rewards enables progress in relatively straightforward tasks such as reach and dial turn. However, it encounters significant challenges in more complex ones, exemplified by the door and hammer tasks within the dexterous hand manipulation domain. The incorporation of pure novelty-seeking rewards (RND) unsurprisingly enhances the RL agent’s exploration, particularly evident in addressing moderately complex tasks like coffee push and assembly. Nevertheless, the performance on dexterous hand manipulation tasks is still unsatisfactory due to the lack of expert-instructed exploration at prohibitively large configuration space. Conversely, AMP exclusively employs expert-guided rewards to incentivize the exploration of expert-like behaviors. While it generally outperforms RND in simpler tasks, its efficacy diminishes in more complex ones such as lever pull and hammer tasks.
Comparison with VIPER. The above observations suggest that the combination of expert-instructed rewards and novelty-seeking rewards is likely to perform favorably. Despite the incorporation of such a combination in the VIPER reward, its empirical performance unexpectedly falls short of both RND and AMP, irrespective of reward standardization. This phenomenon is in line with two limitations outlined in Section 4.2 and further verified in Figure 2, indicating VIPER’s struggle in capturing complex video distributions within intricate tasks. In sharp contrast to VIPER, our proposed reward not only utilizes the generative modeling capabilities of diffusion models, but also adopts conditional entropy as stochastic rewards to bring efficient exploration. The results showcase remarkable performance improvements of 38% and 35% over the best-performing baselines with the same training steps across MetaWorld and Adroit, respectively
@@ 5.3 Zero-Shot Reward Generalization
Figure 9: Ablations. Success rate curves for ablated versions of Diffusion Reward, aggregated over Door and Hammer from Adroit. (a) We test the combinations of generative models and rewards to show the benefit of estimating conditional entropy with the diffusion model. (b) We ablate the choice of the number of denoising steps. (c) We demonstrate that the inherent randomness of Diffusion Reward from the reverse process helps RL exploration. (d) We ablate the choice of reward coefficient. (e) We test the effect of the number of conditional frames. Results are means of 3 seeds with 95% confidence intervals (shaded area). Red is our default.
@@ 6 Conclusion
Appendix
Figure 10: Task descriptions. (left) 10 seen training tasks from MetaWorld and Adroit. (right) 8 unseen tasks from MetaWorld.
@@ Appendix B Task Descriptions
Figure 11: Ablations on diffusion process. The results in (a) suggest that overmuch denoised samples (i.e., 4) may hinder the exploration due to the low variance of estimated entropy. This is further verified in (b) where an appropriate choice of sampling noise scale results in more productive explorations. Results are means of 3 seeds with std. error (shaded area). Red is our default.
@@ Appendix C More Video Prediction Results
Figure 12: Effect of sparse environmental reward. We demonstrate that incorporating sparse environmental reward as a task completion signal is necessary for solving complex manipulation tasks. Results are means of 3 seeds with 95% confidence intervals (shaded area). Red is our default.
@@ Appendix D More Ablations
Sparse environmental reward. The sparse environmental reward $r^{\mathrm{spar}}$ is integrated in Eq. (5), as the environmental supervision of completion is helpful for RL [25]. We remove $r^{\mathrm{spar}}$ to study its effect. The results in Figure 12 show that the performances decrease dramatically without sparse environmental rewards, indicating that the signal of task completion is necessary for solving complex manipulation tasks. Interestingly, more favorable performance is observed in the Door task. We attribute this to the overlap supervision of RND reward and sparse reward, i.e., exploring novel states (door opening) is partially equivalent to completing the task.
@@ Appendix E More Generalization Experiments
Figure 13: More results of reward generalization. Diffusion Reward exhibit better generalization ability than VIPER. Results are means of 3 seeds with 95% confidence intervals (shaded area).

### 18 Mar 2024 v2

@@ 1 Introduction
Figure 1: Overview. (top) We present a reward learning framework in RL using video diffusion models. We perform diffusion processes conditioned on historical frames to estimate conditional entropy as rewards to encourage RL exploration of expert-like behaviors. (bottom) The mean success rate of 10 visual robotic manipulation tasks demonstrates the effectiveness of our Diffusion Reward over 5 runs. Shaded areas are standard errors.
@@ 3 Preliminaries
Real VideoGPT 0.768 (±$\pm 0.141$) 23.35 (±$\pm 4.13$) 0.1017 (±$\pm 0.0670$)
VQ-Diffusion 0.856 (±$\pm 0.114$) 27.46 (±$\pm 5.24$) 0.0596 (±$\pm 0.0450$)
Sim. VideoGPT 0.967 (±$\pm 0.021$) 32.28 (±$\pm 3.15$) 0.0067 (±$\pm 0.0032$)
VQ-Diffusion 0.976 (±$\pm 0.009$) 33.55 (±$\pm 2.31$) 0.0053 (±$\pm 0.0029$)
@@ 4.3 Training Details
Figure 6: Main results. Success rate for our method and baselines on 7 gripper manipulation tasks from MetaWorld and 3 dexterous manipulation tasks from Adroit with image observations. Our method achieves prominent performance on all tasks, and significantly outperforms baselines on complex door and hammer tasks. Results are means of 5 runs with standard errors (shaded area).
@@ 5.1 Experimental Setup
Figure 7: Success rate curves on 5 unseen MetaWorld tasks. Diffusion Reward could generalize to unseen tasks directly and produce reasonable rewards, largely surpassing other baselines. Results are means of 4 runs with standard errors (shaded area).
@@ 5.2 Main Results
Comparison with non-pretraining methods. We present the learning curves of success rates for each method over two simulation domains in Figure 1 and 6. The results show that solely using sparse task rewards enables progress in relatively straightforward tasks such as Reach and Dial Turn. However, it encounters significant challenges in more complex ones, exemplified by the Door and Hammer within the dexterous hand manipulation domain. The incorporation of pure novelty-seeking rewards (RND) unsurprisingly enhances the RL agent’s exploration, particularly evident in addressing moderately complex tasks like Coffee Push and Assembly. Nevertheless, the performance on dexterous hand manipulation tasks is still unsatisfactory due to the lack of expert-instructed exploration at prohibitively large configuration space. Conversely, AMP explicitly employs expert-guided rewards to incentivize the exploration of expert-like behaviors. While it generally outperforms RND in simpler tasks, its efficacy diminishes in more complex ones such as Lever Pull and Hammer.
Comparison with reward pretraining method. The above observations suggest that the combination of expert-instructed rewards and novelty-seeking rewards is likely to perform favorably. Despite the incorporation of such a combination in the VIPER reward, its empirical performance unexpectedly falls short of both RND and AMP. We posit this is attributed to the significantly varied scales of expert-instructed rewards (i.e., log-likelihood), which may eliminate the effect of novelty-seeking rewards. VIPER-std alleviates this issue and shows better performance than the non-standardized version, especially on Coffee Push and Pen, while still underperforming our method. These results are in line with two limitations outlined in Section 4.2 and further verified in Figure 2, indicating VIPER’s struggle in capturing complex video distributions within intricate tasks. In sharp contrast to VIPER, our proposed method not only utilizes the modeling capabilities of diffusion models, but also uses conditional entropy as a reward function to accelerate exploration. The results showcase remarkable performance improvements of 38% and 35% over the best-performing baselines with the same training steps across MetaWorld and Adroit, respectively.
@@ 5.3 Zero-Shot Reward Generalization
Figure 9: Ablations. Success rate curves for ablated versions of Diffusion Reward, aggregated over Door and Hammer from Adroit. (a) We test the combinations of generative models and rewards to show the benefit of estimating conditional entropy with the diffusion model. (b) We ablate the choice of the number of denoising steps. (c) We demonstrate that the inherent randomness of Diffusion Reward from the reverse process helps RL exploration. (d) We ablate the choice of reward coefficient. (e) We test the effect of the number of conditional frames. Results are means of 3 seeds with standard error (shaded area). Red is our default.
@@ Appendix B Task Descriptions
Figure 11: Ablations on diffusion process. The results in (a) suggest that overmuch denoised samples (i.e., 4) may hinder the exploration due to the low variance of estimated entropy. This is further verified in (b) where an appropriate choice of sampling noise scale results in more productive explorations. Results are means of 3 seeds with standard errors (shaded area). Red is our default.
@@ Appendix C More Video Prediction Results
Figure 12: Effect of sparse environmental reward. We demonstrate that incorporating sparse environmental reward as a task completion signal is necessary for solving complex manipulation tasks. Results are means of 3 seeds with standard errors (shaded area).
@@ Appendix D More Ablations
Sparse environmental reward. The sparse environmental reward $r^{\mathrm{spar}}$ is integrated into Eq. (5), as the environmental supervision of completion is helpful for RL. We remove $r^{\mathrm{spar}}$ to study its effect. The results in Figure 12 show that the performances decrease dramatically without sparse environmental rewards, indicating that the signal of task completion is necessary for solving complex manipulation tasks. Interestingly, more favorable performance is observed in the Door task. We attribute this to the overlap supervision of RND reward and sparse reward, i.e., exploring novel states (door opening) is partially equivalent to completing the task.
@@ Appendix E More Generalization Experiments
Figure 13: More results of reward generalization. Diffusion Reward exhibit better generalization ability than VIPER. Results are means of 5 seeds with standard errors (shaded area).